Which of the equations is a line parallel to 4x – 2y = 7?
2y = 4x + 7
–4x – 2y = –7
–2y = 4x + 2
4x + 2y = 2

Answer by richwmiller(17219)

the first one
They both have a slope of two
4x – 2y = 7
4x-7=2y
2x-7/2=y
y = 2x-7/2
–4x – 2y = –7 becomes
y = 2 x+7/2